

12.18, 2007
Day: 33
Weather: Fine but Cloudy
Site: C0002 (NT3-01) (33°18’ N, 136°38’ E)
Science Operations in Labs: Continued physical property and interstitial water chemistry analyses for Hole C0002D cores. Site Summary Meeting for C0002 was held.
Overall Science Update: Completed all science operation. Analytical results and preliminary interpretations for Site C0002 were reported from each specialty group. Four lithostratigraphic units were defined (Unit I-III for Kumano Basin sediments, Unit IV for accretionary prism). Clear differences were observed between Unit IV and other units in terms of structure, chemistry, physical properties, etc.

12.17, 2007
Day: 32
Weather: Fine but Cloudy
Site: C0002 (NT3-01) (33°18’ N, 136°38’ E)
Overall Science Update: Downhole temperature measurement was carried out with APCT-3 on cores 12H, 14H, 16H and 18H, and obtained temperature profile down to 159 mbsf.

12.16, 2007
Day: 31
Weather: Fine but Cloudy
Site: C0002 (NT3-01) (33°18’ N, 136°38’ E)
Science Operations in Labs: Started coring of the riser tophole section at Site C0002. Cores from the Holes C and D are dedicated mainly to geo-engineering assessments for riser BOP deployment. As well as sampling of whole-rounds and shear strength measurements, non-destructive core measurements were carried out as usual. Temperature measurements by APCT3 were successfully conducted in the Hole C0002D. Started writing Site Summary for C0002, in parallel to the remaining discrete sample measurements.

12.15, 2007
Day: 30
Weather: Fine but Cloudy
Site: C0002 (NT3-01) (33°18’ N, 136°38’ E)
Science Operations in Labs: Scientists of expedition 315 and 316 had a meeting for discussing about core processing flow of the next hole (Hole C0002C) and schedule of the site summary report of Site C0002. Continued to complete core processing and measurements.

12.14, 2007
Day: 29
Weather: Fine but Cloudy
Site: C0002 (NT3-01) (33°18’ N, 136°38’ E)
Science Operations in Labs: Drafts of Site Summary reports of Site C0001 were in review by Co-Chiefs and Expedition Project Manager. Continued core processing and measurements. Structural geologists carefully describe deformation structures on split halves of accretionary prism cores.

12.13, 2007
Day: 28
Weather: Overcast
Site: C0002 (NT3-01) (33°17’ N, 136°38’ E)
Science Operations in Labs: Half of the Science Party disembarked. Half of Exp 316 Science Party came embarked instead. Cross-over between two Science Parties is being proceeded through processing of cores from the Hole C0002B. H.
Overall Science Update: Drilled through the base of the Kumano Basin around 920 mbsf, and penetrated into the top of the accretionary prism unit. Lithology and structures were drastically changed.

12.12, 2007
Day: 27
Weather: Cloudy
Site: C0002 (NT3-01) (33°17’ N, 136°38’ E)
Science Operations in Labs: Continued routine core processing on cores from the Hole C0002B. Had a science seminar on recent advances of deepsea exploration for dynamic fluid processes along the edges of the Pacific by Samuel Hulme.

12.11, 2007
Day: 26
Weather: Cloudy
Site: C0002 (NT3-01) (33°17’ N, 136°38’ E)
Science Operations in Labs: Had an End-of-Expedition meeting and discussed about post-expedition studies and host of the second post-expedition meeting. Continued routine core processing on cores from the Hole C0002B. Continued iodine measurement by ICP-MS for the samples from the Site C0001.

12.10, 2007
Day: 25
Weather: Cloudy
Site: C0002 (NT3-01) (33°17’ N, 136°38’ E)
Overall Science Update: Microfossil age determination demonstrated sediments of the Kumano Basin are significantly younger than the sediments of the same depth in the Site C0001.

12.9, 2007
Day: 24
Weather: Fine but Cloudy
Site: C0002 (NT3-01) (33°17’ N, 136°38’ E)
Science Operations in Labs: Completed sample preparation and equipment set up for I analyses in interstitial water (by ICP-MS). Had a science seminar on the results of IODP Exp. 308 Gulf of Mexico by Jan Behrmann (IFM-GEOMAR; DE).

12.8, 2007
Day: 23
Weather: Fine but Cloudy
Site: C0002 (NT3-01) (33°17’ N, 136°38’ E)
Science Operations in Labs: Completed NGR re-measurement on the split halves. Continued inorganic carbon analysis and metal element analyses in interstitial water (by ICP-MS). Had a Site Summary meeting for C0001.
Overall Science Update: Preliminary results of each specialty on the Site C0001 were presented in the Site Summary meeting. Major changes across the boundary at 210 mbsf were recognized in many ways.

12.7, 2007
Day: 22
Weather: Fine but Cloudy
Site: C0001 (NT2-03) (33°14’ N, 136°42’ E)
Science Operations in Labs: Started writing Site Summary reports of C0001. Continued NGR re-measurement for the split halves. Completed XRD analyses.

12.6, 2007
Day: 21
Weather: Fine but Cloudy
Site: C0001 (NT2-03) (33°14’ N, 136°42’ E)
Science Operations in Labs: Core-log-seismic integration specialist found the good correlation between natural gamma ray measurements of core samples and GR data of LWD in the upper sequence. Science party had a seminar on subduction input of the Nankai Trough by Prof. Mike Underwood.

12.5, 2007
Day: 20
Weather: Fine but Cloudy
Site: C0001 (NT2-03) (33°14’ N, 136°42’ E)
Science Operations in Labs: Restarted XRD analysis, which was repaired on 4 Dec by replacement of high-voltage unit. Continued NGR measurements using split halves. Advanced measurements mainly on inorganic chemistry is ongoing. Scientific Party has almost finished the first drafts of the Methods chapter of the Expedition Report, which will be reviewed by Co-Chief Scientists and Expedition Project Manager.

12.4, 2007
Day: 19
Weather: Fine but Cloudy
Site: C0001 (NT2-03) (33°14’ N, 136°42’ E)
Science Operations in Labs: P-wave velocity and electric resistivity measurements for discrete samples are ongoing. Continued NGR measurements on the cores, which had not measured in a coring period due to machine trouble. Corrections on P-wave velocity and electric resistivity by discrete samples measurements using 3rd-party tools enabled better correlation with log data from obtained on Exp. 314.

12.3, 2007
Day: 18
Weather: Fine but Cloudy
Site: C0001 (NT2-03) (33°14’ N, 136°42’ E)
Science Operations in Labs: Finished all personal samplings from split halves including the Hole C0001B, which was taken on the previous Expedition. Re-measured MSCL-W NGR, which could not be done due to mechanical failure. Science seminar was held, and Drs. Tetsuro Hirono (Japan) and Wen-Lu Zhu (USA) gave talks on the summary of the Taiwan Chelungpu-Fault Drilling Project (TCDP).

12.2, 2007
Day: 17
Weather: Fine but Cloudy
Site: C0001 (NT2-03) (33°14’ N, 136°42’ E)
Science Operations in Labs: Science party completed personal samplings from split halves of the Hole C0001E, F and H cores. Micropaleontologists presented preliminary results of biostratigraphy at Holes E and F in cooperation with paleomagnetists.

12.1, 2007
Day: 16
Weather: Fine but Cloudy
Site: C0001 (NT2-03) (33°14’ N, 136°42’ E)
Science Operations in Labs: Science party began writing Expedition Report, mainly on “method” chapter by each specialty group. Continued personal samplings from split halves in non-coring period, which was originally planned during casing operation in the latter half of the expedition. That is because we had so many whole round core samplings during first coring operation, which is a time consuming and complicated process.


11.30, 2007
Day: 15
Weather: Overcast
Site: C0001 (NT2-03) (33°14’ N, 136°42’ E)
Science Operations in Labs: Routine core processing and personal sampling have been conducted. Lab-technician skillfully accomplished trial makings of thin sections for semiconsolidated rock.

11.29, 2007
Day: 14
Weather: Rain
Site: C0001 (NT2-03) (33°14’ N, 136°42’ E)
Science Operations in Labs: No coring operation due to bad weather. Routine core processing and personal sampling from split halves were conducted more smoothly. Lab technicians considerably improved noises on MSCL-image although strong vibrations were still continued due to bad weather. Scientists started to operate some non-routine measurements such as ICP-AES and XRF Core Logger with the assistance of lab technicians. Some scientists became seasick. However large the Chikyu is, ship is a ship.

11.28, 2007
Day: 13
Weather: Rain
Site: C0001 (NT2-03) (33°14’ N, 136°42’ E)
Science Operations in Labs: Science Party continued routine core processing and started individual sampling from split halves. Strong vibration due to bad weather since last evening caused noises on MSCL-image, which should be reprocessed later. Structural geologists are reconstructing paleostress fields using minor fault geometry with paleomagnetic restoration. Co-chiefs and EPM informed about updated operation schedule at the science meeting.

11.27, 2007
Day: 12
Weather: Rain
Site: C0001 (NT2-03) (33°14’ N, 136°42’ E)
Science Operations in Labs: Recovery RCB cores is highly variable but quality is generally good enough to take samples for shore based experimental studies. Science Party continued to measure and described core samples. Physical property specialists estimated the thermal gradient based on downhole data (APCT3) and core thermal conductivity. It is high quality fundamental information about heat flow in Accretionary complex, which also helps temperature estimation for future riser drilling and long-term borehole monitoring. Scan images focusing on shear planes with high (dense) and low (porous) CT-values. Guidance to the Science Party about Expedition-related publication obligation was held, presented by Publication Assistant.

11.26, 2007
Day: 11
Weather: Cloudy
Site: C0001 (NT2-03) (33°14’ N, 136°42’ E)
Science Operations in Labs: As receiving RCB cores from the Hole C0001H, Science Party started core measurements and descriptions. The first two cores were from the duplicated intervals with the ESCS cores at the Hole C0001F, and CT-scan images show less drilling disturbances for RCB cores than ESCS ones. Science party finished thermal conductivity measurement for whole-round cores, and start measurement with discrete cube samples, instead. Structural geologists prepared to make thin sections of healed faults and shear bands with the assistance of a newly arrived lab-technician having special skill on thin section work, and reinvestigated CT-scan images focusing on shear planes with high (dense) and low (porous) CT-values.

11.25, 2007
Day: 10
Weather: Fine but Cloudy
Site: C0001 (NT2-03) (33°14’ N, 136°42’ E)
Science Operations in Labs: Completed first-round descriptions and measurements of the Hole C0001F cores (19 HPCS and 2 ESCS cores). Confirmed sampling procedures from working half cores among the Science Party. In addition to the general agreement, sampling from structurally important intervals were discussed separately in a working group.

11.24, 2007
Day: 9
Weather: Fine but Cloudy
Site: C0001 (NT2-03) (33°14’ N, 136°42’ E)
Science Operations in Labs: Measurements and descriptions of the Hole C0001F cores were continuously conducted. Science seminar on summary of Exp. 314 LWD results was organized and discussed mainly about relationships between log data and expected core-scale deformations.

11.23, 2007
Day: 8
Weather: Fine but Cloudy
サイト:C0001(NT2-03)(33°14’ N, 136°42’ E)
Science Operations in Labs: ESCS coring was conducted two runs due to decrease of penetration by HPCS. Recovery was almost full, but severe drilling disturbance was observed in CT-scan images. Therefore, Co-chiefs and EPM decided to finish ESCS and switch to RCB. Science Party continued to measure and described the cores, and attempted the correlation between Hole E and Hole F. Structural geologists carefully investigated minor normal faults and other deformation structures.

11.22, 2007
Day: 7
Weather: Fine but Cloudy
Site: C0001 (NT2-03) (33°14’ N, 136°42’ E)
Science Operations in Labs: HPCS cores were taken about every 2 hours, and Science Party continued to conduct core processing, descriptions and measurements. Co-chief and EPM gathered information about problems on core processing flow and revised them. Core liner was busted during the cutting operation of 10H Core at the Core Cutting Area probably due to gassy sediments, as well as . Co-chiefs and EPM discuss with Lab Officer about safer procedures.

11.21, 2007
Day: 6
Weather: Fine but Cloudy
Site: C0001 (NT2-03) (33°14’ N, 136°42’ E)
Science Operations in Labs: Co-chiefs and EPM discussed with the drilling operation group about cases of less advance of HPCS coring assembly and had agreement about boundary conditions of finishing HPCS. We also decided cancellation of DVTP operation due to shortage of equipments.
Science Party continued to conduct core processing, descriptions and measurements. The core processing flow is getting faster after making some arrangements on the original plan. Sample Allocation Committee planed sampling party from the working halves during a casing operation except for time critical samples.

11.20, 2007
Day: 5
Weather: Cloudy
Site: C0001 (NT2-03) (33°14’ N, 136°42’ E)
Science Operations in Labs: We have taken 10 HPCS cores about every 2 hours during last 24hrs with satisfactory recovery. We had cracking of core liner for some cores in the Core Cutting Area because of gassy sediments. Science party started to conduct core processing and measurements. Large expansion of sediments made some sections longer than 165 cm, that is the maximum length of MSCR-W. Many whole-round samples were collected for shorebased studies. There was some confusion for whole-round sampling procedures, which caused a stuck of core flow, but it is getting better now. Definition and handling of geotech 'community' samples seemed to be misunderstood by some scientists, which may cause another confusion. Four temperature measurements have successfully been made using APCT 3.

11.19, 2007
Day: 4
Weather: Cloudy
Site: C0001 (NT2-03) (33°14’ N, 136°42’ E)
Science Operations in Labs: Science party continued core processing and measurements of C0001 Hole B cores in the morning, and stopped them to stand ready for arrival of the first core. Paleomagnetists and paleontologists continue to work for statigraphic age determination. Started regular core processing by getting nice 3D X-ray CT scan images. First temperature measurement with APCT3 on the second core, which yielded good data.

11.18, 2007
Day: 3
Weather: Fine but Cloudy
Site: C0001 (NT2-03) (33°14’ N, 136°42’ E)
Science Operations in Labs: Operations to run the drill pipes were stopped in the morning due to failure in the Drawworks motor. Co-Chief Scientists and Expedition Project Manager made a plan for science activities during maintenance of Drawworks. Description and measurement on the Hole B cores taken during the Exp 314, to complement the next coring and to check core flow and pre-spud meeting were carried out. Science meetings about curation, general core flow and measurement plan and lecture for data registration to the J-CORES were held. Science party took 1 hour ship tour under the guide of Safety Manager. Having a full set of Lab Techs onboard, lab preparation for coring is ongoing.

11.17, 2007
Day: 2
Weather: Cloudy
Site: C0001 (NT2-03) (33°14’ N, 136°42’ E)
Science Operations in Labs: Whole science party got onboard. Started 24 hrs working shift after having a safety induction and a short introductory science meeting. Checked measurement procedures and preparation of laboratory instruments by all scientists. Confirmed procedures of downhole temperature measurement with downhole engineer, core tech and OSI.

11.16, 2007
Day:1
Weather: Overcast
Site: C0001 (NT2-03) (33°14’ N, 136°42’ E)
Science Operations in Labs: Lab techs continued set up including dry run of core flow. Onboard Sample Allocation Committee had a discussion about sample and data request and curatorial planning. Had a discussion about coring with core tech.
